OpenStudy (anonymous):

Suppose that D + E = A, that N/A = .25 and that D/A = 0.60. Solve for N/E

OpenStudy (anonymous):

D=.6A

OpenStudy (anonymous):

E=A-(.6A)

OpenStudy (anonymous):

E=.4A A=E/.4

OpenStudy (anonymous):

Thanks I have about ten questions like this and I have no idea how to do them> I am an online student and live in Phoenix but I attend GSU in Atlanta. I haven't been able to get my textbook yet and the Professor begins the semester with this.

OpenStudy (anonymous):

haha sure thing :)

OpenStudy (anonymous):

I don't know how to give medals on here but I will if you tell me how.

OpenStudy (anonymous):

N/A=.25 N=.25A

OpenStudy (anonymous):

click best response

OpenStudy (anonymous):

its next to the metal pic in blue :)

OpenStudy (anonymous):

N=.25(E/.4)

OpenStudy (anonymous):

N/E=.25/.4 N/E= .625 and there is your answer :)

OpenStudy (anonymous):

do you understand my working?

OpenStudy (anonymous):

Sorry I had to answer the door. I think so

OpenStudy (anonymous):

well check it over its just the constant use of simultaneous equations

OpenStudy (anonymous):

also if you want you can become a fan if ur interested

OpenStudy (anonymous):

I did that

OpenStudy (anonymous):

u understand it :)

OpenStudy (anonymous):

not yet but i am going over your work and can probably use it as an example once I get it

OpenStudy (anonymous):

kk message me if anything